# A Text-guided Protein Design Framework

### v2

@@ Abstract
Current AI-assisted protein design mainly utilizes protein sequential and structural information. Meanwhile, there exists tremendous knowledge curated by humans in the text format describing proteins’ high-level functionalities. Yet, whether the incorporation of such text data can help protein design tasks has not been explored. To bridge this gap, we propose ProteinDT, a multi-modal framework that leverages textual descriptions for protein design. ProteinDT consists of three subsequent steps: ProteinCLAP which aligns the representation of two modalities, a facilitator that generates the protein representation from the text modality, and a decoder that creates the protein sequences from the representation. To train ProteinDT, we construct a large dataset, SwissProtCLAP, with 441K text and protein pairs. We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 10 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
@@
To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [101, 6, 102, 8, 9, 10, 11]. Motivated by recent breakthroughs of foundation models [86, 88, 89, 90], approaches in the computational chemistry domain that utilize textual descriptions of drugs in addition to their intrinsic chemical and structural information have proved to be effective in small-molecule drug discovery [16, 95, 96, 97, 138]. This naturally raises a question for the protein domain: how can such a multi-modal mechanism benefit protein engineering and generation? To answer this question, we initiate a novel paradigm that leverages textual descriptions for protein design.
@@
We consider three downstream tasks to verify the versatile functionalities of ProteinDT. The first task is text-to-protein generation, where we follow the ProteinDT pipeline to generate protein sequences from the input text prompts, which describe the desired protein attributes. Using AR and ProteinDiff, the optimal retrieval accuracy can reach over 90%. The second task is zero-shot text-guided protein editing, where the inputs are text prompts and protein sequences. We introduce two editing methods: latent interpolation and latent optimization. Latent interpolation conducts an interpolation in the sequence-level representation space, while latent optimization directly optimizes the token-level representation. Both editing methods inject the text modality and the learned representation is used for protein generation accordingly. We observe that ProteinDT can reach consistently the best hit ratio on 10 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks. Additional qualitative results offer further compelling evidence supporting the validity of ProteinDT. The last task is protein property prediction, aiming to evaluate the robustness and generalization ability of the learned representation from ProteinCLAP. Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain consistently superior performance on four of six benchmark tasks.
Figure 1: Pipeline for ProteinDT framework (a-c) and downstream tasks (d-f). (a) ProteinCLAP, a contrastive learning paradigm, aligns the representation space of the text and protein sequence modalities. (b) ProteinFacilitator model augments the mapping from text sequence representation to protein sequence representation. (c) A protein sequence decoder, which generates protein sequences conditioned on the representations produced from previous steps. (d) Downstream text-to-protein generation task. (e) Downstream text-guided protein design task. (f) Downstream protein property prediction task.
@@ Results
Protein sequence ${\bm{x}}_{\text{p}}$ and representation ${\bm{z}}_{\text{p}}$. Proteins have four levels of structure [25]: primary, secondary, tertiary, and quaternary structure, and the corresponding degree of complexity in the polypeptide chain increases. In this work, we are interested in the protein primary structure ${\bm{x}}_{\text{p}}$, i.e., the amino acid sequence. There are in total 20 different amino acids in proteins, represented with 20 characters, as shown in Figure 1. For the protein sequence representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [26]. It is one of the most advanced Transformer models for modeling sequential data [27]. Additionally, we take the BERT model with ProtBERT pretraining [102]: it was pretrained on a large protein sequence corpus [28, 29], and the pretraining task was to reconstruct the masked tokens.
Text sequence ${\bm{x}}_{\text{t}}$ and representation ${\bm{z}}_{\text{t}}$. The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$. We further adapt the pretrained SciBERT [132], which was pretrained on the full-text computer science and biomedical papers from Semantic Scholar [31]. Notice that although both modalities use BERT-like models, the detailed architectures and token vocabulary are different. Further, such a disentangled design enables us to adopt the respective pretrained models (e.g., ProtBERT and SciBERT) to mitigate the limited availability of structured text-protein sequence pairs for training.
@@
As illustrated in Figure 1, through leveraging the constructed SwissProtCLAP dataset, ProteinDT attains the aforementioned goal with three key steps: a pretraining model, a facilitator, and a decoder. First, the ProteinCLAP model in ProteinDT utilizes contrastive learning to construct representations that integrate knowledge from both text and protein sequences. By doing so, ProteinCLAP not only forms an informative representation of proteins, but also enables the text-protein representation alignment, enabling rich cross-modality tasks like text-protein retrieval. The second step is to train a ProteinFacilitator model, which is an augmented alignment module in addition to ProteinCLAP. To be more concrete, the ProteinFacilitator model maps a piece of text prompt (e.g., “A protein that is stable.") to a corresponding protein representation that captures both the semantic information of the given text and essential protein sequence patterns. Last but not least, ProteinDT subsequently utilizes a conditional generative model to decode the formed protein representation into the protein sequence. We propose one autoregressive model (AR) with Transformer and two diffusion models (ProteinDiff). More details of three pretraining steps can be found in the Methods section. With such a ProteinDT framework, we can explore a broad range of challenging protein design tasks, as will be illustrated below.
@@
Figure 2: Visualization of text-to-protein generation and text-guided protein editing. (a) Visualization for evaluation of text-to-protein generation. The pretrained ProteinCLAP is used to calculate the similarity between the sampled text and generated protein sequence pairs. (b-c) Two methods for text-guided protein editing: latent interpolation and latent optimization. (d-f) Visualization for evaluation of text-guided protein editing. For three types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
@@
Galactica ChatGPT ProteinDT (Ours) w/o ProteinFacilitator w/ ProteinFacilitator AR AR AR ProteinDiff AR ProteinDiff Transformer RNN Transformer Transformer RNN Transformer T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50 T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00 T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
Experiment setting. We randomly sample 200 text sequences from SwissProtCLAP as the text prompts and then generate a protein sequence for each following the ProteinCLAP-ProteinFacilitator-decoder pipeline and ProteinCLAP-decoder pipeline. For evaluation, we adopt the well-trained ProteinCLAP to obtain the retrieval accuracy: each protein sequence is compared to the input text prompt and $T-1$ randomly chosen texts using ProteinCLAP, and a match is successful if the actual text-sequence pair ranks the highest among all comparisons. The evaluation pipeline is shown in Figure 2.
@@
Observations. The main results are presented in Table 1, revealing three key observations. Observation 1. Adopting the ProteinFacilitator model is deemed essential. As shown in Table 1, across all ProteinDT decoder models, incorporating ProteinFacilitator yields higher accuracies compared to models without it. Especially for decoders like AR-Transformer, the performance gaps are substantial by up to around 50%. Observation 2. For the three decoder variants of ProteinDT, AR-Transformer can beat two diffusion models (ProteinDiff). Similar findings are also reported in recent studies on text-to-text generation [35, 36]. We speculate that either the AR model excels in fitting discrete and sequential data models, or we haven’t found the most expressive diffusion model for such a challenging task. These are interesting ML research questions for future exploration. Observation 3. When compared to baseline models, the AR-Transformer with ProteinFacilitator reaches consistently better performance, while the two variants of ProteinDiff with ProteinFacilitator obtain similar performance than Galactica and ChatGPT. It is noteworthy that both baselines are Transformer-based autoregressive models.
@@
Latent interpolation. We have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively. Then we apply the spherical linear interpolation (slerp) [89], which yields an intermediate representation $\tilde{\bm{z}}_{\text{c}}=\text{slerp}({\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}},\theta)$, and the output protein is generated from decoder, as $\tilde{\bm{x}}_{\text{p}}\sim p({\bm{x}}_{\text{p}}|\tilde{\bm{z}}_{\text{c}})$. Here $\theta\in[0,1]$ is a coefficient controlling how close $\tilde{\bm{x}}_{\text{p}}$ is to the text prompt or the input protein, and the larger value of $\theta$ expects $\tilde{\bm{x}}_{\text{p}}$ to be closer to the text prompt.
@@
Latent optimization. We also propose using latent optimization [90, 97] for text-guided protein editing. The high-level idea is to learn a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt. Then, the learned ${\bm{x}}_{w}$ are mapped to the protein sequence using a pretrained decoder $g$, i.e., $\tilde{\bm{x}}_{\text{p}}=g({\bm{z}}_{w})$. Notice that $g$ is a token-wise decoder, and can be trained in an auto-encoder manner with the fixed ProteinCLAP protein encoder. More details of the two methods are in Supplementary C.
@@
Editing tasks and evaluation metrics. To fully verify the effectiveness of ProteinDT, we design 10 zero-shot text-guided protein editing tasks. We further categorize them into three settings as detailed below. For editing evaluation, we employ the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt. The scores are provided by oracle evaluators on each protein sequence, as discussed below. To mitigate bias resulting from different oracle evaluators, we adopt two distinct types of evaluators for each task. Certain evaluations require proteins’ 3D structures, for which we employ the ESMFold algorithm [8]. The pipeline of the task design and evaluation are in Figure 2.
@@
(1) Structure editing. Our initial focus involves four editing tasks associated with the secondary structure [138]. The prompts are in the format of “modify the amino acid sequence to have more (less) alpha helices” and “modify the amino acid sequence to have more (less) beta sheets”. We adopt two methods to measure the secondary structures for each amino acid: (1) A strong protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [37].
@@
(2) Stability editing. Then we consider four types of stability editing tasks. Two proteins are considered [137]: villin HP35 and Pin1 WW domain. The prompts are in the format of “modify the amino acid sequence to have higher (lower) stability”. We consider two evaluators: (1) A strong protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) The predicted local distance difference test (pLDDT) score measuring the confidence for folding, which has been used as a proxy for stability [2].
@@
(3) Peptide binding editing. The last two tasks are peptide binding editing. Peptides are small proteins (usually less than 15 amino acids) with specific functionalities through binding to the target proteins. Here we take 154 pairs of peptide-protein binding pairs from Protein Data Bank (PDB) [39], where the proteins have textual descriptions from UniProt in SwissProtCLAP. Then, the task is to make the input peptide with higher/lower binding affinity with respect to the target protein (textual description). For scoring with the oracle evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [140]; (2) GNINA to calculate the docking score before and after editing [142].
@@
Observations. The main results are in Table 2. We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on all 10 tasks and 20 evaluation methods. However, for the stability editing tasks, latent interpolation methods tend to favor lower stability, while latent optimization can reach a good balance between higher and lower stability tasks. This suggests that latent optimization is in general a preferable method for protein editing. Despite being a generic foundation model, ChatGPT is occasionally the second-best performer on some of the stability editing tasks.
We further provide qualitative results in Figure 3. Figure 3 (a) and (d) show the structures of two input protein sequences. Then ProteinDT adopts the prompts with key words “more $\alpha$-sheets” and “more $\beta$-strands”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e). Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-sheets” and “less $\beta$-strands”. For peptide editing illustration, we take PDB 3IQI, Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example. The input peptide is shown in Figure 3 (g), and the output peptides with higher and lower binding affinity are shown in Figure 3 (h, i). We marked the hydrogen bonds in dashed yellow lines, and there are 6, 7, and 2 bonds in Figure 3 (g, h, i), respectively.
@@
Structure: $\alpha$-helices Structure: $\beta$-sheets Stability: Villin Stability: Pin1 binding editing more less more less more stable less stable more stable less stable higher lower BERT DSSP BERT DSSP BERT DSSP BERT DSSP BERT pLDDT BERT pLDDT BERT pLDDT BERT pLDDT BindingNet GNINA BindingNet GNINA Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 7.81 7.81 73.44 67.19 1.41 2.82 8.45 7.04 26.61 21.77 33.87 25.00 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 29.69 17.19 78.12 18.75 28.17 18.31 39.44 36.62 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 17.19 1.56 82.81 98.44 4.23 4.23 95.77 94.37 58.87 29.84 37.90 70.16 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 9.38 1.56 93.75 98.44 2.82 5.63 97.18 94.37 50.81 29.03 43.55 67.74 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 6.25 1.56 93.75 98.44 1.41 7.04 98.59 92.96 53.23 29.84 45.97 71.77 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 50.00 29.69 54.69 76.56 52.11 42.25 47.89 56.34 39.52 37.90 60.48 61.29
@@
Experiment setting. We consider the following six tasks gathered from TAPE [139]. Secondary structure (SS) prediction [43] is a token-level prediction task, and the goal is to predict the secondary structure (helix, strand, etc.) for each amino acid. SS-Q3 and SS-Q8 refer to the three-way and eight-way classifications, respectively. Homology prediction [44, 45] is to predict the protein fold of the protein sequence. The evaluation metric is accuracy for these three tasks. Contact [46, 47, 48] is to predict if pairs of amino acids within a protein sequence pair are in contact ($<8\accentset{\circ}{A}$ apart), and the metric is the precision for the top $L/2$ contacts, where $L$ is the length of the medium- and long-range proteins. Fluorescence [49] is to predict the log-fluorescence intensity of a protein. Stability [137] is to predict a proxy for protein stability. Both Fluorescence and Stability are regression tasks, and they are measured using Spearman’s $\rho$ metric. More details of the datasets are in Supplementary B.
@@
Observations. The property prediction results are in Table 3. We show two ProtBERT models followed with ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively. We can observe that both ProteinCLAP pretrained representations can consistently reach the best performance on three structure-related prediction tasks. Yet for the other two tasks, Homology and Stability, they are still within the top-three models. The only exception is the Fluorescence task, where all the ProtBERT models are performing worse than other baselines. Noticeably, ProteinCLAP outperforms OntoProtein on five out of six tasks.

### v4

@@ Abstract
Current AI-assisted protein design mainly utilizes protein sequential and structural information. Meanwhile, there exists tremendous knowledge curated by humans in the text format describing proteins’ high-level functionalities. Yet, whether the incorporation of such text data can help protein design tasks has not been explored. To bridge this gap, we propose ProteinDT, a multi-modal framework that leverages textual descriptions for protein design. ProteinDT consists of three subsequent steps: ProteinCLAP which aligns the representation of two modalities, a facilitator that generates the protein representation from the text modality, and a decoder that creates the protein sequences from the representation. To train ProteinDT, we construct a large dataset, SwissProtCLAP, with 441K text and protein pairs. We quantitatively verify the effectiveness of ProteinDT on three challenging tasks: (1) over 90% accuracy for text-guided protein generation; (2) best hit ratio on 12 zero-shot text-guided protein editing tasks; (3) superior performance on four out of six protein property prediction benchmarks.
@@
To this end, most existing such methods merely model the intrinsic properties of protein sequences and structures [5, 6, 7, 8, 9, 10]. Motivated by recent breakthroughs of foundation models [11, 12, 13, 14], approaches in the computational chemistry domain that utilize textual descriptions of drugs in addition to their intrinsic chemical and structural information have proved to be effective in small-molecule drug discovery [15, 16, 17, 18, 19]. This naturally raises a question for the protein domain: how can such a multi-modal mechanism benefit protein engineering and generation? To answer this question, we initiate a novel paradigm that leverages textual descriptions for protein design.
@@
We consider three downstream tasks to verify the versatile functionalities of ProteinDT. The first task is text-to-protein generation, where we follow the ProteinDT pipeline to generate protein sequences from the input text prompts, which describe the desired protein attributes. Using AR and ProteinDiff, the optimal retrieval accuracy can reach over 90%. The second task is zero-shot text-guided protein editing, where the inputs are text prompts and protein sequences. We introduce two editing methods: latent interpolation and latent optimization. Latent interpolation conducts an interpolation in the sequence-level representation space, while latent optimization directly optimizes the token-level representation. Both editing methods inject the text modality and the learned representation is used for protein generation accordingly. We observe that ProteinDT can reach the overall best hit ratio on 12 editing tasks, ranging from structure- and stability-aware tasks, to peptide-binding editing tasks. Additional qualitative results offer further compelling evidence supporting the validity of ProteinDT. The last task is protein property prediction, aiming to evaluate the robustness and generalization ability of the learned representation from ProteinCLAP. Compared to six state-of-the-art protein sequence representation methods, ProteinDT can obtain superior performance on four of six benchmark tasks.
Figure 1: Pipeline for ProteinDT pretraining framework (a-c) and downstream tasks (d-f). (a) ProteinCLAP, a contrastive learning paradigm, aligns the representation space of the text and protein sequence modalities. (b) ProteinFacilitator model augments the mapping from text sequence representation to protein sequence representation. (c) A protein sequence decoder, which generates protein sequences conditioned on the representations produced from previous steps. (d) Downstream text-to-protein generation task. (e) Downstream text-guided protein editing task. (f) Downstream protein property prediction task.
@@ Results
Protein sequence ${\bm{x}}_{\text{p}}$ and representation ${\bm{z}}_{\text{p}}$. Proteins have four levels of structure [24]: primary, secondary, tertiary, and quaternary structure, and the corresponding degree of complexity in the polypeptide chain increases. In this work, we are interested in the protein primary structure ${\bm{x}}_{\text{p}}$, i.e., the amino acid sequence. There are in total 20 different amino acids in proteins, represented with 20 characters, as shown in Figure 1. For the protein sequence-level representation ${\bm{z}}_{\text{p}}=f_{\text{p}}({\bm{x}}_{\text{p}})$, we use BERT as the encoder [25]. It is one of the most advanced Transformer models for modeling sequential data [26]. Additionally, we take the BERT model with ProtBERT pretraining [6]: it was pretrained on a large protein sequence corpus [27, 28], and the pretraining task was to reconstruct the masked tokens.
Text sequence ${\bm{x}}_{\text{t}}$ and representation ${\bm{z}}_{\text{t}}$. The textual descriptions ${\bm{x}}_{\text{t}}$ are formulated in the form of sequences of tokens, and we also adopt the BERT model as the encoder to get the sequence-level representation ${\bm{z}}_{\text{t}}=f_{\text{t}}({\bm{x}}_{\text{t}})$. We further adapt the pretrained SciBERT [29], which was pretrained on the full-text computer science and biomedical papers from Semantic Scholar [30]. Notice that although both modalities use BERT-like models, the detailed architectures and token vocabulary are different. Further, such a disentangled design enables us to adopt the respective pretrained models (e.g., ProtBERT and SciBERT) to mitigate the limited availability of structured text-protein sequence pairs for training.
@@
As illustrated in Figure 1, through leveraging the constructed SwissProtCLAP dataset, ProteinDT attains the aforementioned goal with three key steps: a pretraining model, a facilitator, and a decoder. First, the ProteinCLAP model in ProteinDT utilizes contrastive learning to construct representations that integrate knowledge from both text and protein sequences. By doing so, ProteinCLAP not only forms an informative representation of proteins, but also enables the text-protein representation alignment, enabling rich cross-modality tasks like text-protein retrieval. The second step is to train a ProteinFacilitator model, which is an augmented alignment module in addition to ProteinCLAP. To be more concrete, the ProteinFacilitator model maps a piece of text prompt (e.g., “A protein that is stable.") to a corresponding protein representation that captures both the semantic information of the given text and essential protein sequence patterns. Last but not least, ProteinDT subsequently utilizes a conditional generative model to decode the formed protein representation into the protein sequence. We propose one autoregressive model (AR) and two diffusion models (ProteinDiff). More details of the three pretraining steps can be found in the Methods section.
@@
Figure 2: Visualization of text-to-protein generation and text-guided protein editing. (a) Visualization for evaluation of text-to-protein generation. The pretrained ProteinCLAP is used to calculate the similarity between the sampled text and generated protein sequence pairs. (b-c) Two methods for text-guided protein editing: latent interpolation and latent optimization. (d-e) Visualization for evaluation of text-guided protein editing. For four types of editing tasks, different evaluation metrics (marked in red) are applied accordingly.
@@
Galactica ChatGPT ProteinDT (Ours)
w/o ProteinFacilitator w/ ProteinFacilitator
AR AR AR ProteinDiff AR ProteinDiff
Transformer RNN Transformer Transformer RNN Transformer
T = 4 51.5 38.5 49.00 24.00 35.50 97.00 40.50 51.50
T = 10 29.0 23.0 27.00 10.50 17.50 91.00 21.50 25.00
T = 20 19.0 15.5 20.00 5.50 9.50 83.50 15.00 13.50
Experiment setting. We randomly sample 200 text sequences from SwissProtCLAP as the text prompts and then generate a protein sequence for each following the ProteinCLAP-ProteinFacilitator-decoder pipeline and ProteinCLAP-decoder pipeline. For evaluation, we adopt the well-trained ProteinCLAP to obtain the retrieval accuracy: each protein sequence is compared to the input text prompt and $T-1$ randomly chosen texts using ProteinCLAP, and a match is successful if the actual text-sequence pair ranks the highest among all comparisons. The evaluation pipeline is shown in Figure 2, and please check Supplementary D for more details.
@@
Observations. The main results are presented in Table 1, revealing three key observations. Observation 1. Adopting the ProteinFacilitator model is deemed essential. As shown in Table 1, across all ProteinDT decoder models, incorporating ProteinFacilitator yields higher accuracies compared to models without it. Especially for decoders like AR-Transformer, the performance gaps are substantial by up to around 50%. Observation 2. For the three decoder variants of ProteinDT, AR-Transformer can beat the two diffusion models (ProteinDiff). Similar findings are also reported in recent studies on text-to-text generation [34, 35]. We speculate that either the AR model excels in fitting discrete and sequential data models, or we haven’t found the most expressive diffusion model for such a challenging task. These are interesting ML research questions for future exploration. Observation 3. When compared to baseline models, the AR-Transformer with ProteinFacilitator reaches consistently better performance, while the two variants of ProteinDiff with ProteinFacilitator obtain similar performance than Galactica and ChatGPT. It is noteworthy that both baselines are Transformer-based autoregressive models.
@@
Latent interpolation. Latent interpolation aims at getting the latent codes by directly interpolating two endpoint representations [12, 13, 36], where the latent codes will be used for sampling data points through the decoder. In this setting, we have two representations for the text prompt and input protein as ${\bm{z}}_{\text{t}}$ and ${\bm{z}}_{\text{p}}$, respectively. Then we apply the spherical linear interpolation (SLERP) [13], which yields an intermediate representation $\tilde{\bm{z}}_{\text{c}}=\text{SLERP}({\bm{z}}_{\text{t}},{\bm{z}}_{\text{p}},\theta)$, and the output protein is generated from decoder, as $\tilde{\bm{x}}_{\text{p}}\sim p({\bm{x}}_{\text{p}}|\tilde{\bm{z}}_{\text{c}})$. Here $\theta\in[0,1]$ is a coefficient controlling how close $\tilde{\bm{x}}_{\text{p}}$ is to the text prompt or the input protein, and the larger value of $\theta$ expects $\tilde{\bm{x}}_{\text{p}}$ to be closer to the text prompt.
@@
Latent optimization. We also propose using latent optimization [14, 37, 18] for text-guided protein editing. The high-level idea is to directly optimize a latent code in the representation space and then decode it to get the sampled data. Specifically in ProteinDT, it learns a token-level latent code ${\bm{z}}_{w}$ that is close to both the text and protein sequence-level representation, as ${\bm{z}}_{w}=\arg\min(\lambda\|P({\bm{z}}_{w})-{\bm{z}}_{\text{t}}\|^{2}+(1-\lambda)\|P({\bm{z}}_{w})-{\bm{z}}_{\text{p}}\|^{2})$, where $P$ is the pooling function, and $\lambda$ controls the closeness to the target prompt. Then, the learned ${\bm{x}}_{w}$ are mapped to the protein sequence using a pretrained decoder $g$, i.e., $\tilde{\bm{x}}_{\text{p}}=g({\bm{z}}_{w})$. Notice that $g$ is a token-wise decoder, and can be trained in an auto-encoder manner with the fixed ProteinCLAP protein encoder. More details of the two methods are in Supplementary C.
@@
Editing tasks and evaluation metrics. To fully verify the effectiveness of ProteinDT, we design 12 zero-shot text-guided protein editing tasks. We further categorize them into four settings as detailed below. For editing evaluation, we employ the satisfactory hit ratio (%), which measures the ratio of the output score being higher or lower than the input score, as indicated in the text prompt. The scores are provided by evaluators on each protein sequence, as discussed below. To mitigate bias resulting from different evaluators, we adopt two distinct types of evaluators for each task (except region editing). Certain evaluations require proteins’ 3D structures, for which we employ the ESMFold algorithm [7]. The pipeline of the task design and evaluation are in Figure 2.
@@
(1) Structure editing. Our initial focus involves four editing tasks associated with the secondary structure [19]. The prompts are in the format of “modify the amino acid sequence to have more (less) alpha helices” and “modify the amino acid sequence to have more (less) beta sheets”. We adopt two methods to measure the secondary structures for each amino acid: (1) A protein sequence evaluator for secondary structure prediction (as will be discussed in the property prediction downstream below); (2) Define Secondary Structure of Proteins (DSSP) algorithm for assigning secondary structure to the amino acids [38].
(2) Region editing. The second editing task is region editing. The predicted Local Distance Difference Test (pLDDT) is a per-residue confidence metric used to estimate the reliability of the predicted 3D structure of a protein sequence. We consider proteins from this paper [39], which reveals the relation between ordered/disordered regions and pLDDT. Ordered regions most frequently show higher pLDDT, while the disordered regions are reflected with low pLDDT. The prompts are in the format of “modify the amino acid sequence to have more ordered/disordered regions”. We adopt the pLDDT from ESMFold as the evaluation metric [7].
@@
(3) Stability editing. Then we consider four types of thermodynamic stability editing tasks. Two proteins are considered [40]: villin HP35 and Pin1 WW domain. The prompts are in the format of “modify the amino acid sequence to have higher (lower) stability”. We consider two evaluators: (1) A protein sequence evaluator for stability (as will be discussed in the property prediction downstream below); (2) Rosetta to quantify the energy after structure prediction, in Rosetta Energy Units (REU) [41, 42]. The score function is REF2015 [43].
@@
(4) Peptide binding editing. The last two tasks are peptide binding editing. Peptides are small proteins (usually less than 15 amino acids) with specific functionalities through binding to the target proteins. Here we take 154 pairs of peptide-protein binding pairs from Protein Data Bank (PDB) [44], where the proteins have textual descriptions from UniProt in SwissProtCLAP. Then, the task is to make the input peptide have higher/lower binding affinity with respect to the target protein (textual description). For scoring with the evaluator, we consider two options: (1) A pretrained deep learning model for binding affinity prediction [45]; (2) GNINA to calculate the docking score before and after editing [46].
@@
Observations. The main results are in Table 2. We can observe that two variants of ProteinDT (latent interpolation and latent optimization) generally perform better on 12 tasks and 21 evaluation methods. However, for the stability editing tasks, latent interpolation methods tend to favor lower stability, while latent optimization can reach a good balance between higher and lower stability tasks. This suggests that latent optimization is, in general, a preferable method for protein editing. Despite being a generic foundation model, ChatGPT is occasionally the second-best performer on some of the stability editing tasks.
We further provide qualitative results in Figure 3. Figure 3 (a) and (d) show the structures of two input protein sequences. Then ProteinDT adopts the prompts with keywords “more $\alpha$-helices” and “more $\beta$-sheets”, the edited protein structures (edited sequences folded with ESMFold) are shown in Figure 3 (b) and (e). Similarly for Figure 3 (c) and (f) when using prompts like “less $\alpha$-helices” and “less $\beta$-sheets”. For peptide editing illustration, we take PDB 3IQI [47], Haemophilus influenzae O-acetylserine sulfhydrylase in complex with the peptide MNENI, as an example. The input peptide is shown in Figure 3 (g), and the output peptides with higher and lower binding affinity are shown in Figure 3 (h, i). We marked the hydrogen bonds in dashed yellow lines, and there are 6, 7, and 2 bonds in Figure 3 (g, h, i), respectively.
@@
Structure: $\alpha$-helices Structure: $\beta$-sheets Region more less more less ordered disordered BERT DSSP BERT DSSP BERT DSSP BERT DSSP pLDDT pLDDT Galactica 5.46 7.02 13.26 10.53 1.95 2.14 11.50 11.31 8.47 23.73 ChatGPT 5.46 6.82 7.80 6.43 4.68 2.92 5.07 3.51 6.78 62.71 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 67.64 75.63 19.49 10.14 27.88 11.31 53.61 69.98 5.08 94.92 ProteinDiff-BERT 38.79 51.27 48.15 34.70 39.38 15.79 42.30 65.50 3.39 94.92 AR-BERT 27.49 45.61 56.92 39.38 9.16 5.26 69.79 73.10 15.25 84.75 Latent Optimization 51.66 54.19 33.92 32.94 46.59 39.57 39.38 46.59 6.78 91.53 Stability: Villin Stability: Pin1 binding editing more stable less stable more stable less stable higher lower BERT Rosetta BERT Rosetta BERT Rosetta BERT Rosetta BindingNet GNINA BindingNet GNINA Galactica 7.81 51.56 73.44 31.25 1.41 4.23 8.45 4.23 26.61 21.77 33.87 25.00 ChatGPT 29.69 57.81 78.12 28.12 28.17 19.72 39.44 71.83 16.94 12.90 26.61 35.48 ProteinDT (Ours) Latent Interpolation ProteinDiff-RNN 17.19 50.00 82.81 54.69 4.23 26.76 95.77 70.42 58.87 29.84 37.90 70.16 ProteinDiff-BERT 9.38 23.44 93.75 68.75 2.82 16.90 97.18 84.51 50.81 29.03 43.55 67.74 AR-BERT 6.25 40.62 93.75 59.38 1.41 25.35 98.59 73.24 53.23 29.84 45.97 71.77 Latent Optimization 50.00 53.12 54.69 40.62 52.11 35.21 47.89 63.38 39.52 37.90 60.48 61.29
@@
Experiment setting. We consider the following six tasks gathered from TAPE [48]. Secondary structure (SS) prediction [49] is a token-level prediction task, and the goal is to predict the secondary structure (helix, sheet, etc.) for each amino acid. SS-Q3 and SS-Q8 refer to the three-way and eight-way classifications, respectively. Homology prediction [50, 51] is to predict the protein fold of the protein sequence. The evaluation metric is accuracy for these three tasks. Contact [52, 53, 44] is to predict if pairs of amino acids within a protein sequence pair are in contact ($<8\accentset{\circ}{A}$ apart), and the metric is the precision for the top $L/2$ contacts, where $L$ is the length of the medium- and long-range proteins. Fluorescence [54] is to predict the log-fluorescence intensity of a protein. Stability [40] is to predict a proxy for protein stability. Both tasks are measured using Spearman’s $\rho$ metric.
@@
Observations. The property prediction results are in Table 3. We show two variants of ProteinCLAP pretraining, corresponding to the EBM-NCE loss and InfoNCE loss, respectively. We can observe that both ProteinCLAP pretrained representations can consistently reach the best performance on three structure-related prediction tasks. Yet for the other two tasks, Homology and Stability, they are still within the top three models. The only exception is the Fluorescence task, where all the ProtBERT models are performing worse than other baselines.